\begin{thebibliography}{30}
\addcontentsline{toc}{section}{References}




\bibitem{B}
A. Bejancu,
\newblock {\em Geometry of CR-submaniflods},
\newblock Kluwer Academic, 1986.





\bibitem{BL}
J. P. Bourguignon, H. B. Lawson,
\newblock {\em A mathematician's visit to Kaluza-Klein theory},
\newblock Rend. Semin. Mat. Torino Fasc. Spec. (1989), 143-163.




\bibitem{BL2}
J. P. Bourguignon, H. B. Lawson,
\newblock {\em Stability and isolation phenomena for Yang-mills fields},
\newblock Commum. Math. Phys. (1981), \textbf{79}, 189-230.



\bibitem{BW}
P. Baird, J. C. Wood,
\newblock {\em Harmonic morphisms between Riemannian manifolds},
\newblock Oxford science publications, 2003.




\bibitem{C}
B. Y. Chen,
\newblock {\em Geometry of slant submaniflods},
\newblock Katholieke Universiteit Leuven, Leuven, 1990.




\bibitem{CMMS}
V. Cort\'{e}s, C. Mayer, T. Mohaupt, F. Saueressig,
\newblock {\em Special geometry of Euclidean supersymmetry 1. Vector multiplets},
\newblock J. High Energy Phys. (2004), \textbf{03}, 028.



\bibitem{FIP}
    M. Falcitelli, S. Ianus, and A. M. Pastore,
    \newblock {\em Riemannian submersions and related topics},
    \newblock World Scientific Publishing Co., 2004.





\bibitem{G}
A. Gray,
\newblock {\em Pseudo-Riemannian almost product manifolds and submersions},
\newblock J. Math. Mech, (1967), \textbf{16}, 715-737.





\bibitem{IIMV}
S. Ianus, A. M. Ionescu, R. Mazzocco, G. E. Vilcu,
\newblock {\em Riemannian submersions from almost contact metric manifolds },
\newblock arXiv: 1102.1570v1 [math. DG].





\bibitem{IMV}
S. Ianus, R. Mazzocco, G. E. Vilcu,
\newblock {\em Riemannian submersions from quaternionic manifolds },
\newblock Acta. Appl. Math.(2008) \textbf{104}, 83-89.






\bibitem{IV}
S. Ianus, M. Visinescu,
\newblock {\em Kaluza-Klein theory with scalar fields and generalized Hopf manifolds},
\newblock Class. Quantum Gravity (1987), \textbf{4}, 1317-1325.




\bibitem{IV2}
S. Ianus, M. Visinescu,
\newblock {\em Space-time compactification and Riemannian submersions},
\newblock In: Rassias, G.(ed.) The Mathematical Heritage of C. F. Gauss,
 (1991), 358-371, World Scientific, River Edge.




\bibitem{M}
M. T. Mustafa,
\newblock {\em Applications of harmonic morphisms to gravity},
\newblock J. Math. Phys. (2000), \textbf{41}(10), 6918-6929.





\bibitem{O}
B. O'Neill,
\newblock {\em The fundamental equations of a submersion},
\newblock Mich. Math. J., (1966), \textbf{13}, 458-469.




\bibitem{P}
K. S. Park,
\newblock {\em H-slant submersions},
\newblock Bull. Korean Math. Soc. (2012), \textbf{49} No. 2, 329-338.




\bibitem{P2}
K. S. Park,
\newblock {\em H-semi-invariant submersions},
\newblock Taiwan. J. Math., Accepted.



\bibitem{P3}
K. S. Park,
\newblock {\em H-semi-slant submersions},
\newblock Preprint (2011).


\bibitem{PP}
K. S. Park, R. Prasad
\newblock {\em Semi-slant submersions},
\newblock Preprint (2011).



\bibitem{S}
B. Sahin,
\newblock {\em slant submersions from almost Hermitian manifolds},
\newblock Bull. Math. Soc. Sci. Math. Roumanie Tome (2011), \textbf{54}(102) No. 1,
93 - 105.




\bibitem{S2}
B. Sahin,
\newblock {\em Semi-invariant submersions from almost Hermitian manifolds},
\newblock Canad. Math. Bull. (2011), \textbf{54}, No. 3.



\bibitem{W}
B. Watson,
\newblock {\em Almost Hermitian submersions},
\newblock J. Differential Geom.(1976), \textbf{11}(1), 147-165.





\bibitem{W2}
B. Watson,
\newblock {\em $G,G'$-Riemannian submersions and nonlinear gauge field equations of general relativity},
\newblock In: Rassias, T. (ed.) Global Analysis - Analysis on manifolds,
dedicated M. Morse. Teubner-Texte Math., (1983), \textbf{57},
324-349, Teubner, Leipzig.








\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb,amsmath,amsthm}
\usepackage[dvips]{graphicx}

\let\oldmarginpar\marginpar
\renewcommand\marginpar[1]{\-\oldmarginpar[\raggedleft\footnotesize #1]%
{\raggedright\footnotesize #1}}

\begin{document}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\theoremstyle{definition}
\newtheorem{example}[theorem]{Example}

\def\R{{\mathbb R}}
\def\H{{\mathbb H}}
\def\rank{{\text{rank}\,}}
\def\bd{{\partial}}

\title{V-Semi-slant submersions}

\author{Kwang-Soon Park}
\address{Department of Mathematical Sciences, Seoul National University, Seoul 151-747, Republic of Korea}
\email{parkksn@gmail.com}
%\thanks{The research of the first author was supported by the National Research Foundation of Korea (NRF)
%grant funded by the Korea government (MEST)(No. 2009-0057445).}

%\author{Suhyoung Choi \thanks{S. Choi was supported by the National Research Foundation
%of Korea (NRF) grant funded by the Korea government (MEST)  (No. 2009-0057445).} \and
%Craig D. Hodgson\thanks{C.D. Hodgson was partially supported by Australian Research Council
%grant DP0663399.}\and Gye-Seon Lee\thanks{G.-S. Lee was supported
%by Basic Science Research Program through the National Research
%Foundation of Korea (NRF) funded by the Ministry of Education,Science and
%Technology (KRF-2008-621-C00003).}

%\address{S. Choi \at Department of Mathematical Sciences, KAIST, Daejeon 305-701, Republic of Korea
%\and
%C.D. Hodgson \at Department of Mathematics and Statistics, University of Melbourne, Victoria 3010, Australia
%\and
%G.-S. Lee (\,\Letter\,) \at Department of Mathematical Sciences, KAIST, Daejeon 305-701, Republic of Korea
%}

%\\\email{shchoixk@gmail.com}
%\\\email{craigdh@unimelb.edu.au}
%\\\email{smileabacus@kaist.ac.kr}

\keywords{Riemannian submersion \and slant angle \and totally geodesic}

\subjclass[2000]{53C15; 53C43.}   %Primary 57M50; Secondary 57N16, 53A20, 53C15}

\begin{abstract}
Let $F$ be a Riemannian submersion from an almost Hermitian
manifold $(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. We
introduce the notion of the v-semi-slant submersion. And then we
obtain some properties on it. In particular, we give some examples
for it.
\end{abstract}

\maketitle
\section{Introduction}\label{intro}
\addcontentsline{toc}{section}{Introduction}

Let $F$ be a $C^{\infty}$-submersion from a semi-Riemannian
manifold $(M,g_M)$ onto a semi-Riemannian manifold $(N,g_N)$. Then
according to the conditions on the map $F : (M,g_M) \mapsto
(N,g_N)$, we have the following submersions:

semi-Riemannian submersion and Lorentzian submersion \cite{FIP},
Riemannian submersion (\cite{G}, \cite{O}), slant submersion
(\cite{C}, \cite{S}), almost Hermitian submersion \cite{W},
contact-complex submersion \cite{IIMV}, quaternionic submersion
\cite{IMV}, almost h-slant submersion and h-slant submersion
\cite{P}, semi-invariant submersion \cite{S2}, almost h-semi-invariant
submersion and h-semi-invariant
submersion \cite{P2}, semi-slant submersions \cite{PP}, almost h-semi-slant submersions and h-semi-slant submersions \cite{P3}, etc.
As we know, Riemannian submersions are
related with physics and have their applications in the Yang-Mills
theory (\cite{BL2}, \cite{W2}), Kaluza-Klein theory (\cite{BL},
\cite{IV}), Supergravity and superstring theories (\cite{IV2},
\cite{M}), etc.

The paper is organized as follows. In section 2 we remind some notions which is neeed for this paper.
 In section 3 we give the definition of the v-semi-slant submersion and obtain
some interesting properties on it. In section 4 we construct
some examples of the v-semi-slant submersion.

\section{Preliminaries}\label{prelim}

Let $(M, g_M)$ and $(N, g_N)$ be Riemannian
manifolds, where $M$, $N$ are $C^{\infty}-$manifolds and $g_M$, $g_N$ are Riemannian metrics,  
and $F : M \mapsto N$ a $C^{\infty}-$submersion. The map
$F$ is said to be {\em Riemannian submersion} if the differential
$F_*$ preserves the lengths of horizontal vectors \cite{IMV}. 

Let $(M, g_M, J)$ be an almost Hermitian manifold, where $J$ is an almost complex structure. A Riemannian
submersion $F : (M, g_M, J)\mapsto (N, g_N)$ is called a {\em
slant submersion} if the angle $\theta=\theta(X)$ between $JX$ and the
space $\ker (F_*)_p$ is constant for any nonzero $X\in T_p M$ and
$p\in M$ \cite{S}. 

We call $\theta$ a {\em slant angle}. 

A Riemannian submersion $F : (M, g_M, J)\mapsto (N, g_N)$ is called
a {\em semi-invariant submersion} if there is a distribution
$\mathcal{D}_1 \subset \ker F_*$ such that
$$
\ker F_*=\mathcal{D}_1\oplus \mathcal{D}_2, \
 J(\mathcal{D}_1)=\mathcal{D}_1, \ J(\mathcal{D}_2)\subset (\ker
F_*)^{\perp},
$$
where $\mathcal{D}_2$ is the orthogonal complement of
$\mathcal{D}_1$ in $\ker F_*$ \cite{S}.  

A Riemannian submersion $F : (M,g_M,J)\mapsto
(N,g_N)$ is called a {\em semi-slant submersion} if there is a
distribution $\mathcal{D}_1\subset \ker F_*$ such that
$$
\ker F_* =\mathcal{D}_1\oplus \mathcal{D}_2, \
J(\mathcal{D}_1)=\mathcal{D}_1,
$$
and the angle $\theta=\theta(X)$ between $JX$ and the space
$(\mathcal{D}_2)_p$ is constant for nonzero $X\in
(\mathcal{D}_2)_p$ and $p\in M$, where $\mathcal{D}_2$ is the
orthogonal complement of $\mathcal{D}_1$ in $\ker F_*$. 

We call the angle $\theta$ a {\em semi-slant angle}.

Let $(M, g_M)$ and $(N,
g_N)$ be Riemannian manifolds and $F : (M, g_M) \mapsto (N, g_N)$
a smooth map. The second fundamental form of $F$ is given by
$$
(\nabla F_*)(X,Y) := \nabla^F _X F_* Y-F_* (\nabla _XY) \quad
\text{for} \ X,Y\in \Gamma(TM),
$$
where $\nabla^F$ is the pullback connection and we denote
conveniently by $\nabla$ the Levi-Civita connections of the
metrics $g_M$ and $g_N$ \cite{BW}. Recall that $F$ is said to be {\em
harmonic} if $trace (\nabla F_*)=0$ and $F$ is called a {\em
totally geodesic} map if $(\nabla F_*)(X,Y)=0$ for $X,Y\in \Gamma
(TM)$ \cite{BW}.

\section{V-Semi-slant submersions}\label{semi}

\begin{definition}
Let $(M,g_M,J)$ be an almost Hermitian manifold and $(N,g_N)$ a
Riemannian manifold. A Riemannian submersion $F : (M,g_M,J)\mapsto
(N,g_N)$ is called a {\em v-semi-slant submersion} if there is a
distribution $\mathcal{D}_1\subset (\ker F_*)^{\perp}$ such that
$$
(\ker F_*)^{\perp} =\mathcal{D}_1\oplus \mathcal{D}_2, \
J(\mathcal{D}_1)=\mathcal{D}_1,
$$
and the angle $\theta=\theta(X)$ between $JX$ and the space
$(\mathcal{D}_2)_p$ is constant for nonzero $X\in
(\mathcal{D}_2)_p$ and $p\in M$, where $\mathcal{D}_2$ is the
orthogonal complement of $\mathcal{D}_1$ in $(\ker F_*)^{\perp}$.
\end{definition}

We call the angle $\theta$ a {\em v-semi-slant angle}.

\begin{remark}
Let $F$ be a v-semi-slant submersion from an almost Hermitian manifold $(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then there is a
distribution $\mathcal{D}_1\subset (\ker F_*)^{\perp}$ such that
$$
(\ker F_*)^{\perp} =\mathcal{D}_1\oplus \mathcal{D}_2, \
J(\mathcal{D}_1)=\mathcal{D}_1,
$$
and the angle $\theta=\theta(X)$ between $JX$ and the space
$(\mathcal{D}_2)_p$ is constant for nonzero $X\in
(\mathcal{D}_2)_p$ and $p\in M$, where $\mathcal{D}_2$ is the
orthogonal complement of $\mathcal{D}_1$ in $(\ker F_*)^{\perp}$.

If $\mathcal{D}_2=(\ker F_*)^{\perp}$, then we call the map $F$ a {\em v-slant submersion} and the angle $\theta$ {\em v-slant angle} \cite{S}.
Otherwise, if $\theta=\frac{\pi}{2}$, then we call the map $F$ a {\em v-semi-invariant submersion} \cite{S2}.
\end{remark}

Let $F : (M,g_M,J)\mapsto (N,g_N)$ be a v-semi-slant
submersion. Then there is a distribution $\mathcal{D}_1\subset
(\ker F_*)^{\perp}$ such that
$$
(\ker F_*)^{\perp} =\mathcal{D}_1\oplus \mathcal{D}_2, \
J(\mathcal{D}_1)=\mathcal{D}_1,
$$
and the angle $\theta=\theta(X)$ between $JX$ and the space
$(\mathcal{D}_2)_p$ is constant for nonzero $X\in
(\mathcal{D}_2)_p$ and $p\in M$, where $\mathcal{D}_2$ is the
orthogonal complement of $\mathcal{D}_1$ in $(\ker F_*)^{\perp}$.

Then for $X\in \Gamma((\ker F_*)^{\perp})$, we have
$$
X = PX+QX,
$$
where $PX\in \Gamma(\mathcal{D}_1)$ and $QX\in
\Gamma(\mathcal{D}_2)$.

For $X\in \Gamma(\ker F_*)$, we get
$$
JX = \phi X+\omega X,
$$
where $\phi X\in \Gamma(\ker F_*)$ and $\omega X\in \Gamma((\ker
F_*)^{\perp})$.

For $Z\in \Gamma((\ker F_*)^{\perp})$, we obtain
$$
JZ = BZ+CZ,
$$
where $BZ\in \Gamma(\ker F_*)$ and $CZ\in \Gamma((\ker
F_*)^{\perp})$.

For $U\in \Gamma(TM)$, we have
$$
U = \mathcal{V}U+\mathcal{H}U,
$$
where $\mathcal{V}U\in \Gamma(\ker F_*)$ and $\mathcal{H}U\in
\Gamma((\ker F_*)^{\perp})$.

Then
$$
\ker F_* = B \mathcal{D}_2 \oplus \mu,
$$
where $\mu$ is the orthogonal complement of $B \mathcal{D}_2$
in $\ker F_*$ and is invariant  under $J$. Furthermore,

{\setlength\arraycolsep{2pt}
\begin{eqnarray*}
& & C \mathcal{D}_1 = \mathcal{D}_1, B \mathcal{D}_1 = 0, C \mathcal{D}_2 \subset \mathcal{D}_2,
\omega(\ker F_*) = \mathcal{D}_2        \\
& & \phi^2+B\omega = -id, C^2+\omega B = -id, \omega \phi +C\omega
= 0, BC+\phi B = 0.
\end{eqnarray*}}

Define the tensors $\mathcal{T}$ and $\mathcal{A}$ by
\begin{align*}
  &\mathcal{A}_E F= \mathcal{H}\nabla_{\mathcal{H}E} \mathcal{V}F+\mathcal{V}\nabla_{\mathcal{H}E} \mathcal{H}F \\
   &\mathcal{T}_E F= \mathcal{H}\nabla_{\mathcal{V}E} \mathcal{V}F+\mathcal{V}\nabla_{\mathcal{V}E}
   \mathcal{H}F
\end{align*}
for vector fields $E, F$ on $M$, where $\nabla$ is the Levi-Civita
connection of $g_M$. Define
$$
\widehat{\nabla}_X Y := \mathcal{V}\nabla_X Y \quad \text{for} \ X,Y\in \Gamma(\ker F_*).
$$
Then we easily have

\begin{lemma}
Let $(M,g_M,J)$ be a K\"{a}hler manifold and $(N,g_N)$ a
Riemannian manifold. Let $F : (M,g_M,J) \mapsto (N,g_N)$ be a
v-semi-slant submersion. Then we get
\begin{enumerate}
\item
\begin{align*}
  &\widehat{\nabla}_X \phi Y+\mathcal{T}_X \omega Y = \phi\widehat{\nabla}_X Y+B\mathcal{T}_X Y    \\
  &\mathcal{T}_X \phi Y+\mathcal{H}\nabla_X \omega Y =
  \omega\widehat{\nabla}_X Y+C\mathcal{T}_X Y
\end{align*}
for $X,Y\in \Gamma(\ker F_*)$.
\item
\begin{align*}
  &\mathcal{V}\nabla_Z BW+\mathcal{A}_Z CW = \phi\mathcal{A}_Z W+B\mathcal{H}\nabla_Z W    \\
  &\mathcal{A}_Z BW+\mathcal{H}\nabla_Z CW = \omega\mathcal{A}_Z
  W+C\mathcal{H}\nabla_Z W
\end{align*}
for $Z,W\in \Gamma((\ker F_*)^{\perp})$.
\item
\begin{align*}
  &\widehat{\nabla}_X BZ+\mathcal{T}_X CZ = \phi\mathcal{T}_X Z+B\mathcal{H}\nabla_X Z    \\
  &\mathcal{T}_X BZ+\mathcal{H}\nabla_X CZ =
  \omega\mathcal{T}_X Z+C\mathcal{H}\nabla_X Z
\end{align*}
for $X\in \Gamma(\ker F_*)$ and $Z\in \Gamma((\ker F_*)^{\perp})$.
\end{enumerate}
\end{lemma}

\begin{theorem}
Let $F$ be a v-semi-slant submersion from an almost Hermitian
manifold $(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then
the slant distribution $\mathcal{D}_2$ is integrable if and only
if we obtain
$$
\mathcal{A}_X Y = 0 \quad \text{and} \quad PC(\nabla_X Y-\nabla_Y X)=0
$$
for $X,Y\in \Gamma(\mathcal{D}_2)$
\end{theorem}

\begin{proof}
For $X,Y\in \Gamma(\mathcal{D}_2)$ and $Z\in \Gamma(\mathcal{D}_1)$, assume that $\mathcal{A}_X Y=\frac{1}{2}\mathcal{V}[X,Y] = 0$, we obtain
\begin{align*}
g_M ([X,Y], JZ)&= -g_M (J(\nabla_X Y-\nabla_Y X), Z)    \\
      &= -g_M (B\nabla_X Y+C\nabla_X Y-B\nabla_Y X-C\nabla_Y X, Z)    \\
      &= -g_M (C(\nabla_X Y-\nabla_Y X), Z).
\end{align*}
Therefore, we have the result.
\end{proof}

Similarly, we get

\begin{theorem}
Let $F$ be a v-semi-slant submersion from an almost Hermitian
manifold $(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then
the complex distribution $\mathcal{D}_1$ is integrable if and only
if we have
$$
\mathcal{A}_X Y = 0 \quad \text{and} \quad B(\nabla_X Y-\nabla_Y X)=0
$$
for $X,Y\in \Gamma(\mathcal{D}_1)$.
\end{theorem}

\begin{lemma}
Let $(M,g_M,J)$ be a K\"{a}hler manifold and $(N,g_N)$ a Riemannian
manifold. Let $F : (M,g_M,J) \mapsto (N,g_N)$ be a v-semi-slant
submersion. Then the complex distribution $\mathcal{D}_1$ is
integrable if and only if we get
$$
\mathcal{A}_X Y = 0 \quad \text{for} \ X,Y\in \Gamma(\mathcal{D}_1). 
$$
\end{lemma}

\begin{proof}
For $X,Y\in \Gamma(\mathcal{D}_1)$ and $Z\in \Gamma(\ker F_*)$, assume that $\mathcal{A}_X Y = 0$, we have
\begin{align*}
g_M ([X,Y], \omega Z)&=g_M ([X,Y], JZ)= -g_M (J(\nabla_X Y-\nabla_Y X), Z)    \\
      &= -g_M (\mathcal{A}_X JY+\mathcal{H}\nabla_X JY-\mathcal{A}_Y JX -\mathcal{H}\nabla_Y JX, Z)        \\
      &= -g_M (\mathcal{A}_X JY-\mathcal{A}_Y JX, Z).
\end{align*}
Therefore, the result follows.
\end{proof}

In a similar way, we have

\begin{lemma}
Let $(M,g_M,J)$ be a K\"{a}hler manifold and $(N,g_N)$ a
Riemannian manifold. Let $F : (M,g_M,J) \mapsto (N,g_N)$ be a
v-semi-slant submersion. Then the slant distribution $\mathcal{D}_2$
is integrable if and only if we obtain
$$
\mathcal{A}_X Y = 0 \quad \text{and} \quad P((\mathcal{A}_X BY-\mathcal{A}_Y BX)+ \mathcal{H}(\nabla_X CY-\nabla_Y CX))=0
$$
for $X,Y\in \Gamma(\mathcal{D}_2)$.
\end{lemma}

\begin{proposition}\label{prop:slant}
Let $F$ be a v-semi-slant submersion from an almost Hermitian manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then we obtain
$$
C^2 X = -\cos^2 \theta X \quad \text{for} \ X\in
\Gamma(\mathcal{D}_2),
$$
where $\theta$ denotes the v-semi-slant angle of $\mathcal{D}_2$.
\end{proposition}

\begin{proof}
Since
$$
\cos \theta = \frac{g_M (JX, C X)}{|JX|\cdot|C X|} =
\frac{-g_M (X, C^2 X)}{|X|\cdot|C X|}
$$
and $\displaystyle{ \cos \theta = \frac{|C X|}{|JX|} }$, we have
$$
\cos^2 \theta = -\frac{g_M (X, C^2 X)}{|X|^2} \quad \text{for} \
X\in \Gamma(\mathcal{D}_2).
$$
Hence,
$$
C^2 X = -\cos^2 \theta X \quad \text{for} \ X\in
\Gamma(\mathcal{D}_2).
$$
\end{proof}

\begin{remark}
It is easy to see that the converse of Proposition
\ref{prop:slant} is also true.
\end{remark}

Assume that the v-semi-slant angle $\theta$ is not equal to
$\displaystyle{\frac{\pi}{2}}$ and define an endomorphism
$\widehat{J}$ of $(\ker F_*)^{\perp}$ by
$$
\widehat{J} := JP+\frac{1}{\cos \theta}CQ.
$$
Then,
\begin{eqnarray} \label{compst}
{\widehat{J}}^2 = -id \quad \text{on} \ (\ker F_*)^{\perp}.
\end{eqnarray}

From (\ref{compst}), we have

\begin{theorem}
Let $F$ be a v-semi-slant submersion from an almost Hermitian manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$ with the v-semi-slant
angle $\theta\in [0,\frac{\pi}{2})$. Then $N$ is an even-dimensional
manifold.
\end{theorem}

\begin{proposition}
Let $F$ be a  v-semi-slant submersion from a K\"{a}hler manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then the
distribution $\ker F_*$ defines a totally geodesic foliation if and
only if
$$
\omega (\widehat{\nabla}_X \phi Y+\mathcal{T}_X \omega
Y)+C(\mathcal{T}_X \phi Y+\mathcal{H}\nabla_X \omega Y) = 0 \quad
\text{for} \ X,Y\in \Gamma(\ker F_*).
$$
\end{proposition}

\begin{proof}
For $X,Y\in \Gamma(\ker F_*)$,
\begin{align*}
\nabla_X Y&= -J\nabla_X JY= -J(\widehat{\nabla}_X \phi Y+\mathcal{T}_X \phi Y+\mathcal{T}_X \omega Y
             +\mathcal{H}\nabla_X \omega Y)   \\
          &= -(\phi \widehat{\nabla}_X \phi Y+\omega\widehat{\nabla}_X \phi Y+B\mathcal{T}_X \phi Y
             +C\mathcal{T}_X \phi Y+\phi \mathcal{T}_X \omega Y+\omega \mathcal{T}_X \omega Y    \\
          & \ \ \  +B\mathcal{H}\nabla_X \omega Y+C\mathcal{H}\nabla_X \omega
             Y).
\end{align*}
Thus,
$$
\nabla_X Y\in \Gamma(\ker F_*) \Leftrightarrow
\omega(\widehat{\nabla}_X \phi Y+\mathcal{T}_X \omega
Y)+C(\mathcal{T}_X \phi Y+\mathcal{H}\nabla_X \omega Y) =0.
$$
\end{proof}

Similarly, we have

\begin{proposition}
Let $F$ be a  v-semi-slant submersion from a K\"{a}hler manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then the
distribution $(\ker F_*)^{\perp}$ defines a totally geodesic
foliation if and only if
$$
\phi(\mathcal{V}{\nabla}_X BY+\mathcal{A}_X CY)+B(\mathcal{A}_X
BY+\mathcal{H}\nabla_X CY) = 0 \quad \text{for} \ X,Y\in
\Gamma((\ker F_*)^{\perp}).
$$
\end{proposition}

\begin{proposition}
Let $F$ be a  v-semi-slant submersion from a K\"{a}hler manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then the
distribution $\mathcal{D}_1$ defines a totally geodesic foliation if
and only if
$$
\phi\mathcal{A}_X JY+B\mathcal{H}\nabla_X JY = 0 \
\text{and} \ Q(\omega\mathcal{A}_X JY+C\mathcal{H}\nabla_X JY) =
0
$$
for $X,Y\in \Gamma(\mathcal{D}_1)$.
\end{proposition}

\begin{proof}
For $X,Y\in \Gamma(\mathcal{D}_1)$, we get
\begin{align*}
\nabla_X Y&= -J\nabla_X JY= -J(\mathcal{A}_X JY+\mathcal{H}\nabla_X JY)   \\
          &= -(\phi\mathcal{A}_X JY+\omega\mathcal{A}_X JY+B\mathcal{H}\nabla_X JY+C\mathcal{H}\nabla_X JY).
\end{align*}
Hence,
$$
\nabla_X Y\in \Gamma(\mathcal{D}_1) \Leftrightarrow \phi\mathcal{A}_X JY+B\mathcal{H}\nabla_X JY = 0 \
\text{and} \ Q(\omega\mathcal{A}_X JY+C\mathcal{H}\nabla_X JY) =
0.
$$
\end{proof}

In a similar way, we obtain

\begin{proposition}
Let $F$ be a  v-semi-slant submersion from a K\"{a}hler manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then the
distribution $\mathcal{D}_2$ defines a totally geodesic foliation if
and only if
\begin{align*}
  &\phi(\mathcal{V}\nabla_X BY+\mathcal{A}_X CY)+B(\mathcal{A}_X BY+\mathcal{H}\nabla_X CY) = 0     \\
  &P(\omega(\mathcal{V}\nabla_X BY+\mathcal{A}_X CY)+C(\mathcal{A}_X BY+\mathcal{H}\nabla_X CY)) = 0
\end{align*}
for $X,Y\in \Gamma(\mathcal{D}_2)$.
\end{proposition}

\begin{theorem}
Let $F$ be a  v-semi-slant submersion from a K\"{a}hler manifold
$(M,g_M,J)$ onto a Riemannian manifold $(N,g_N)$. Then $F$ is a
totally geodesic map if and only if
\begin{align*}
  &\omega(\widehat{\nabla}_X \phi Y+\mathcal{T}_X \omega Y)+C(\mathcal{T}_X \phi Y
    +\mathcal{H}\nabla_X \omega Y) = 0   \\
  &\omega(\widehat{\nabla}_X BZ+\mathcal{T}_X CZ)+C(\mathcal{T}_X BZ
    +\mathcal{H}\nabla_X CZ) = 0
\end{align*}
for $X,Y\in \Gamma(\ker F_*)$ and $Z\in \Gamma((\ker F_*)^{\perp})$.
\end{theorem}

\begin{proof}
Since $F$ is a Riemannian submersion, we obtain
$$
(\nabla F_*)(Z_1,Z_2)=0 \quad \text{for} \ Z_1,Z_2\in \Gamma((\ker
F_*)^{\perp}).
$$
For $X,Y\in \Gamma(\ker F_*)$, we have
\begin{align*}
(\nabla F_*)(X,Y)&= -F_* (\nabla_X Y) = F_* (J\nabla_X (\phi Y+\omega Y))   \\
          &= F_* (\phi \widehat{\nabla}_X \phi Y+\omega\widehat{\nabla}_X \phi Y+B\mathcal{T}_X \phi Y
             +C\mathcal{T}_X \phi Y+\phi \mathcal{T}_X \omega Y+\omega \mathcal{T}_X \omega Y   \\
          & \ \ \  +B\mathcal{H}\nabla_X \omega Y+C\mathcal{H}\nabla_X \omega Y).
\end{align*}
Thus,
$$
(\nabla F_*)(X,Y) = 0 \Leftrightarrow \omega(\widehat{\nabla}_X \phi
Y+\mathcal{T}_X \omega Y)+C(\mathcal{T}_X \phi Y
    +\mathcal{H}\nabla_X \omega Y) = 0.
$$
For $X\in \Gamma(\ker F_*)$ and $Z\in \Gamma((\ker F_*)^{\perp})$,
 we get
\begin{align*}
(\nabla F_*)(X,Z)&= -F_* (\nabla_X Z) = F_* (J\nabla_X (BZ+CZ))   \\
          &= F_* (\phi \widehat{\nabla}_X BZ+\omega\widehat{\nabla}_X BZ+B\mathcal{T}_X BZ
             +C\mathcal{T}_X BZ+\phi \mathcal{T}_X CZ+\omega \mathcal{T}_X CZ   \\
          & \ \ \  +B\mathcal{H}\nabla_X CZ+C\mathcal{H}\nabla_X CZ).
\end{align*}
Hence,
$$
(\nabla F_*)(X,Z) = 0 \Leftrightarrow \omega(\widehat{\nabla}_X
BZ+\mathcal{T}_X CZ)+C(\mathcal{T}_X BZ
    +\mathcal{H}\nabla_X CZ) = 0.
$$
Since $(\nabla F_*)(X,Z)=(\nabla F_*)(Z,X)$, the result follows.
\end{proof}

Let $F : (M,g_M)\mapsto (N,g_N)$ be a Riemannian submersion. The map
$F$ is called a Riemannian submersion {\em with totally umbilical
fibers} if
\begin{eqnarray}\label{umbilic}
\mathcal{T}_X Y = g_M (X, Y)H \quad \text{for} \ X,Y\in \Gamma(\ker
F_*),
\end{eqnarray}
where $H$ is the mean curvature vector field of the fiber.

Then we obtain

\begin{lemma}
Let $F$ be a v-semi-slant submersion with totally umbilical fibers
from a K\"{a}hler manifold $(M,g_M,J)$ onto a Riemannian manifold
$(N,g_N)$. Then we have
$$
H\in \Gamma(\mathcal{D}_2).
$$
\end{lemma}

\begin{proof}
For $X,Y\in \Gamma(\mu)$ and $W\in \Gamma(\mathcal{D}_1)$, we get
$$
\mathcal{T}_X JY+\widehat{\nabla}_X JY = \nabla_X JY = J\nabla_X Y =
B\mathcal{T}_X Y+C\mathcal{T}_X Y+\phi\widehat{\nabla}_X
Y+\omega\widehat{\nabla}_X Y.
$$
Using (\ref{umbilic}), we easily obtain
$$
g_M (X, JY) g_M (H, W) = -g_M (X, Y) g_M (H, JW).
$$
Interchanging the role of $X$ and $Y$, we get
$$
g_M (Y, JX) g_M (H, W) = -g_M (Y, X) g_M (H, JW)
$$
so that combining the above two equations, we have
$$
g_M (X, Y) g_M (H, JW) = 0,
$$
which means $H\in \Gamma(\mathcal{D}_2)$.
\end{proof}

\begin{corollary}
Let $F$ be a v-semi-slant submersion with totally umbilical fibers
from a K\"{a}hler manifold $(M,g_M,J)$ onto a Riemannian manifold
$(N,g_N)$ such that $\mathcal{D}_1=(\ker F_*)^{\perp}$. Then each fiber is minimal.
\end{corollary}

\begin{remark}
Let $F$ be a v-semi-slant submersion from a K\"{a}hler manifold $(M,
g_M, J)$ onto a Riemannian manifold $(N, g_N)$. Then there is a
distribution $\mathcal{D}_1\subset (\ker F_*)^{\perp}$ such that
$$
(\ker F_*)^{\perp} =\mathcal{D}_1\oplus \mathcal{D}_2, \
J(\mathcal{D}_1)=\mathcal{D}_1,
$$
and the angle $\theta=\theta(X)$ between $JX$ and the space
$(\mathcal{D}_2)_p$ is constant for nonzero $X\in (\mathcal{D}_2)_p$
and $p\in M$, where $\mathcal{D}_2$ is the orthogonal complement of
$\mathcal{D}_1$ in $(\ker F_*)^{\perp}$. Furthermore,
$$
C \mathcal{D}_2 \subset \mathcal{D}_2, \quad B \mathcal{D}_2
\subset \ker F_*, \quad \ker F_*=B\mathcal{D}_2\oplus \mu,
$$
where $\mu$ is the orthogonal complement of $B \mathcal{D}_2$
in $\ker F_*$ and is invariant  under $J$. For the curvature tensor, it is sufficient to calculate the
holomorphic sectional curvatures in a K\"{a}hler manifold.

Given a plane $P$ being invariant by $J$ in $T_p M$, $p\in M$, there
is an orthonormal basis $\{ X, JX \}$ of $P$. Denote by $K(P)$,
$K_*(P)$, and $\widehat{K}(P)$ the sectional curvatures of the plane
$P$ in $M$, $N$, and the fiber $F^{-1}(F(p))$, respectively, where
$K_*(P)$ denotes the sectional curvature of the plane $P_* = < F_*X,
F_*JX > $ in $N$. Let $K(X\wedge Y)$ be the sectional curvature of
the plane spanned by the tangent vectors $X, Y\in T_p M$, $p\in M$.
Using both Corollary 1 of [14, p.465] and (1.28) of [7, p.13], we
obtain the following :

\begin{enumerate}

\item If $P\subset (\mu)_p$, then with some computations
we have
$$
K(P)=\widehat{K}(P)+|\mathcal{T}_X X|^2-|\mathcal{T}_X JX|^2-g_M
(\mathcal{T}_X X, J[JX, X]).
$$

\item If $P\subset (\mathcal{D}_2\oplus B\mathcal{D}_2)_p$ with $X\in
(\mathcal{D}_2)_p$, then we get

\begin{align*}
K(P)&= \sin^2\theta \cdot K(X\wedge B X)+2(g_M ((\nabla_X
\mathcal{A})(X,CX), BX)+g_M (\mathcal{A}_X CX, \mathcal{T}_{BX} X)   \\
    & \ \ \ -g_M (\mathcal{A}_{CX} X, \mathcal{T}_{BX} X)-g_M (\mathcal{A}_{X} X, \mathcal{T}_{BX} CX))  +\cos^2\theta \cdot K(X\wedge CX).
\end{align*}

\item If $P\subset (\mathcal{D}_1)_p$, then we obtain
$$
K(P)= K_* (P)-3|\mathcal{V}J\nabla_X X|^2.
$$
\end{enumerate}
\end{remark}

\section{Examples}\label{exam}

\begin{example}
Let $(M,g_M,J)$ be an almost Hermitian manifold. Let $\pi : TM \mapsto M$ be the natural projection.
Then the map $\pi$ is a v-semi-slant submersion such that $\mathcal{D}_1=(\ker \pi_*)^{\perp}$ \cite{FIP}.
\end{example}

\begin{example}
Let $(M,g_M,J)$ be a $2m$-dimensional almost Hermitian manifold and $(N,g_N)$ a $(2m-1)$-dimensional Riemannian manifold.
Let $F$ be a Riemannian submersion from an almost Hermitian manifold $(M,
g_M, J)$ onto a Riemannian manifold $(N, g_N)$. Then the map $F$ is a v-semi-slant submersion such that
$$
\mathcal{D}_1=((\ker F_*)\oplus J(\ker F_*))^{\perp} \quad \text{and} \quad \mathcal{D}_2=J(\ker F_*)
$$
with the v-semi-slant angle $\theta=\frac{\pi}{2}$.
\end{example}

\begin{example}
Define a map $F : \mathbb{R}^6 \mapsto \mathbb{R}^4$ by
$$
F(x_1,x_2,\cdots, x_6) = (x_1,x_3 \sin \alpha-x_5 \cos \alpha,x_6,x_2),
$$
where $\alpha\in (0,\frac{\pi}{2})$. Then the map $F$ is a
v-semi-slant submersion such that
$$
\mathcal{D}_1 = <\frac{\partial}{\partial x_1},
\frac{\partial}{\partial x_2}> \ \text{and} \ \mathcal{D}_2 =
<\frac{\partial}{\partial x_6}, \sin \alpha \frac{\partial}{\partial
x_3}-\cos \alpha \frac{\partial}{\partial x_5}>
$$
with the v-semi-slant angle $\theta=\alpha$.
\end{example}

\begin{example}
Define a map $F : \mathbb{R}^8 \mapsto \mathbb{R}^4$ by
$$
F(x_1,x_2,\cdots, x_8) = (x_4,x_3,\frac{x_5-x_8}{\sqrt{2}},x_6).
$$
Then the map $F$ is a v-semi-slant submersion such that
$$
\mathcal{D}_1 = <\frac{\partial}{\partial
x_3},\frac{\partial}{\partial x_4}> \ \text{and} \ \mathcal{D}_2 =
<\frac{\partial}{\partial x_6}, \frac{\partial}{\partial
x_5}-\frac{\partial}{\partial x_8}>
$$
with the v-semi-slant angle $\theta=\frac{\pi}{4}$.
\end{example}

\begin{example}
Define a map $F : \mathbb{R}^{12} \mapsto \mathbb{R}^5$ by
$$
F(x_1,x_2,\cdots, x_{12}) =
(x_2,\frac{x_5+x_6}{\sqrt{2}},\frac{x_7+x_9}{\sqrt{2}},\frac{x_8+x_{10}}{\sqrt{2}},x_1).
$$
Then the map $F$ is a v-semi-slant submersion such that
$$
\mathcal{D}_1 = <\frac{\partial}{\partial x_1},
\frac{\partial}{\partial x_2},\frac{\partial}{\partial
x_7}+\frac{\partial}{\partial x_9},\frac{\partial}{\partial
x_8}+\frac{\partial}{\partial x_{10}}> \ \text{and} \ \mathcal{D}_2
= <\frac{\partial}{\partial x_5}+\frac{\partial}{\partial x_6}>
$$
with the v-semi-slant angle $\theta=\frac{\pi}{2}$.
\end{example}

\begin{example}
Define a map $F : \mathbb{R}^{10} \mapsto \mathbb{R}^6$ by
$$
F(x_1,x_2,\cdots, x_{10}) =
(\frac{x_3-x_5}{\sqrt{2}},x_6,\frac{x_7+x_9}{\sqrt{2}},x_8,x_1,x_2).
$$
Then the map $F$ is a v-semi-slant submersion such that
$$
\mathcal{D}_1 = <\frac{\partial}{\partial x_1},
\frac{\partial}{\partial x_2}> \ \text{and} \ \mathcal{D}_2 =
<\frac{\partial}{\partial x_6},\frac{\partial}{\partial x_8},\frac{\partial}{\partial x_3}-\frac{\partial}{\partial x_5},
\frac{\partial}{\partial x_7}+\frac{\partial}{\partial
x_9}>
$$
with the v-semi-slant angle $\theta=\frac{\pi}{4}$.
\end{example}

\begin{example}
Define a map $F : \mathbb{R}^8 \mapsto \mathbb{R}^4$ by
$$
F(x_1,x_2,\cdots, x_8) = (x_1,x_3 \cos \alpha-x_5 \sin
\alpha,x_2,x_4\sin \beta+x_6\cos \beta),
$$
where $\alpha$ and $\beta$ are constant. Then the map $F$ is a
v-semi-slant submersion such that
$$
\mathcal{D}_1 = <\frac{\partial}{\partial
x_1},\frac{\partial}{\partial x_2}> \ \text{and} \ \mathcal{D}_2 =
<\cos \alpha\frac{\partial}{\partial x_3}-\sin
\alpha\frac{\partial}{\partial x_5}, \sin
\beta\frac{\partial}{\partial x_4}+\cos
\beta\frac{\partial}{\partial x_6}>
$$
with the v-semi-slant angle $\theta$ with $\cos\theta=|\sin
(\alpha-\beta)|$.
\end{example}

\end{document}